The chance that a student will clear the quiz-1 paper is 0.59 and the chance that he will clear both quiz-1 and quiz-2 papers is 0.17. The chance that he will clear at least one quiz paper is 0.74. What is the chance that he will clear quiz-2 paper?

1
Expert's answer
2022-02-15T11:31:28-0500
"P(1\cup 2)=P(1)+P(2)-P(1\cap 2)"

Then


"P(2)=P(1\cup 2)+P(1\cap 2)-P(1)"

Substitute


"P(2)=0.74+0.17-0.59=0.32"

The chance that a student will clear quiz-2 paper is 0.32.